OpenStudy (anonymous):

An object is dropped from rest and falls freely 20. meters to Earth. When is the speed of the object 9.8 meters per second?

OpenStudy (anonymous):

1) during the entire first second of its fall (2) at the end of its first second of fall (3) during its entire time of fall (4) after it has fallen 9.8 meters

OpenStudy (anonymous):

at the end of its first second of fall

OpenStudy (anonymous):

v = u + at -9.8 = -9.8t t = 1 sec

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

v final velocity u initial velocity a acceleration t time free fall a = -9.8 v is negative as it is going down

OpenStudy (jamesj):

Alternatively: acceleration is by definition change in velocity. Here the acceleration is is a constant, g in the downward direction. Hence of the object starts at rest and has velocity v in the downward direction at time t, then by definition \[ acceleration = \frac{v}{t} = g \] i.e., \( v = gt \) Hence \[ v = g \] when \( t = 1 \) second.